# Microphase Segregation in Polyelectrolyte Brushes

### v2

@@ 0.1 Molecular Origin of Multi-layer Morphologies
We emphasize that the microphase segregation in multi-layer PE brushes is significantly different from that in the pearl-necklace structure of single PE chains in poor solvents. The multi-layer brushes are not simply a parallel alignment of pearl-necklaces. The segregation is not formed by different portions of the same chain, but by different subpopulations of all chains. This arrangement of different polymers in different condensed layers optimizes the total elastic energy caused by chain deformation.
@@
Commonly used characterization techniques like ellipsometry and AFM can only provide the overall brush height, which is not sufficient to effectively reflect the microstructure of multi-layer morphologies in PE brushes. On the other hand, reflectivity is a scattering technique that detects the detailed structure of buried layers by utilizing the interference of X-rays or neutrons. However, extracting density profiles directly from reflectivity spectra is quite challenging because the scattering intensity is measured in reciprocal space 63. The reflected intensity $R$ of an X-ray beam is related to the gradient of the electron density distribution $\mathrm{d}\rho_{e}(z)/\mathrm{dz}$ 64:
$R(Q_{z})/R_{F}=\left|\frac{1}{\rho_{e}^{\infty}}\int\textrm{dz}\ \frac{\mathrm{d}\rho_{e}(z)}{\mathrm{dz}}\exp{(iQ_{z}z)}\right|^{2}~{},$ (6)
@@
Fig. 6 shows the evolution of the reflectivity spectra for different multi-layer structures formed in PE brushes. The key features of the reflectivity spectra are the oscillation period, which is related to layer thickness, and the oscillation amplitude, which is related to the sharpness of the polymer-solvent interface. The spectra exhibit uniform oscillations for morphologies with a single condensed layer ($\chi=2.00$). As the number of condensed layers further increases to two ($\chi=1.70$), three ($\chi=1.45$) and four ($\chi=1.30$), the number of polymer-solvent interfaces that can reflect X-rays also increases, reducing the period of the oscillation. In addition, the irregular distances between these interfaces lead to interference between their reflected X-rays, thus generating complex, non-uniform spectra. The oscillations do not maintain a constant amplitude and period. Furthermore, with the increase of structural heterogeneity, more energy is required for the X-rays to reach the substrate. This increases the momentum $Q_{z}$ required to obtain appreciable amplitude in the reflectivity spectra, and thus shifts the total spectra to higher $Q_{z}$. The addition of the outer diffuse layer in the pre-swollen morphology ($\chi=1.20$) adds slightly more interference to the spectra and reduces the amplitude of the oscillations. Finally, for fully swollen brushes ($\chi=1.00$), the polymer-solvent interface is diffuse and the reflectivity spectra is essentially flat and nearly featureless. Our SCFT calculations thus bridge the experimentally measurable reflectivity spectra to the microstructure of multi-layer brushes.
@@ Conclusions
In the current work, we use a simple system of PE brushes composed of homopolymers with uniform backbone charge density to illustrate the effect of competing interactions on the structural heterogeneity. Our theory can be straightforwardly generalized to brushes with various chain architectures, copolymer compositions and charge patterns, as well as be easily implemented to study protein brushes with specific amino acid sequences. Here, we focus on the multi-layer morphologies formed by microphase segregation of PE brushes in the normal direction. However, solving SCFT in high dimensions for brushes with varying grafting densities may lead to segregation in other directions with rich surface patterns. Furthermore, current applications of PE brushes largely rely on their steric and electrostatic repulsion to protect the surface underneath. As revealed in this work, designing structurally heterogeneous PE brushes by tuning their electrostatic and hydrophobic interactions could regulate surface properties at a higher level.

### v3

@@ 0.1 Molecular Origin of Multi-layer Morphologies
We emphasize that the microphase segregation in multi-layer PE brushes is significantly different from that in the pearl-necklace structure of single PE chains in poor solvents 56, 57, 60. The multi-layer brushes are not simply a parallel alignment of pearl-necklaces. The segregation is not formed by different portions of the same chain, but by different subpopulations of all chains. This arrangement of different polymers in different condensed layers optimizes the total elastic energy caused by chain deformation.
@@
Commonly used characterization techniques like ellipsometry and AFM can only provide the overall brush height, which is not sufficient to effectively reflect the microstructure of multi-layer morphologies in PE brushes. On the other hand, reflectivity is a scattering technique that detects the detailed structure of buried layers by utilizing the interference of X-rays or neutrons. However, extracting density profiles directly from reflectivity spectra is quite challenging because the scattering intensity is measured in reciprocal space 78. The reflected intensity $R$ of an X-ray beam is related to the gradient of the electron density distribution $\mathrm{d}\rho_{e}(z)/\mathrm{dz}$ 79:
$R(Q_{z})/R_{F}=\left|\frac{1}{\rho_{e}^{\infty}}\int\textrm{dz}\ \frac{\mathrm{d}\rho_{e}(\mathrm{z})}{\mathrm{dz}}\exp{(iQ_{z}\mathrm{z})}\right|^{2}~{},$ (6)
@@
Fig. 7 shows the evolution of the reflectivity spectra for different multi-layer structures formed in PE brushes. The key features of the reflectivity spectra are the oscillation period, which is related to layer thickness, and the oscillation amplitude, which is related to the sharpness of the polymer-solvent interface. The spectra exhibit uniform oscillations for morphologies with a single condensed layer ($\chi=2.00$). As the number of condensed layers further increases to two ($\chi=1.70$), three ($\chi=1.45$) and four ($\chi=1.30$), the number of polymer-solvent interfaces that can reflect X-rays also increases. This reduces the period of the oscillations and increases the sharpness of the peaks. In addition, the irregular distances between these interfaces lead to interference between their reflected X-rays, thus generating complex, non-uniform spectra. The oscillations do not maintain a constant amplitude and period. Furthermore, with the increase of structural heterogeneity, more energy is required for the X-rays to reach the substrate. This increases the momentum $Q_{z}$ required to obtain appreciable amplitude in the reflectivity spectra, and thus shifts the total spectra to higher $Q_{z}$. While keeping the same multi-layer morphology, adding more salt increases the screening effect which makes the layers more compact. The resulting spectra also shifts to higher $Q_{z}$ (see Fig. S4 in the SI).
The addition of the outer diffuse layer in the pre-swollen morphology ($\chi=1.20$) adds slightly more interference to the spectra and reduces the amplitude of the oscillations. Finally, for fully swollen brushes ($\chi=1.00$), the polymer-solvent interface is diffuse and the reflectivity spectra is essentially flat and nearly featureless. Therefore, the features of the non-uniform oscillating spectra signify the multi-layer morphologies of PE brushes. The key features of the spectra predicted by our theory are also consistent with those observed in experiments. For instance, Dura et al. found multilayer morphologies in Nafion films which show oscillations in the spectra with a high degree of interference for $Q_{z}>0.06~{}\textup{\AA}^{-1}$ 27. This is close to the onset of non-uniform oscillations found in our theoretical work. In addition, as the number of condensed layers increases, the period of the oscillations were found to increase 28, which is also in agreement with the features shown in Fig. 7.
Our SCFT calculations bridge the experimentally measurable reflectivity spectra to the microstructure of multi-layer brushes. For an experimentally measured spectrum, a series of candidate density profiles can be generated by our SCFT, which can then be converted to corresponding reflectivity spectra. The one with the best agreement with the experimental result can be used to identify the brush morphology.
@@ Conclusions
In the current work, we use a simple system of PE brushes composed of homopolymers with uniform backbone charge density to illustrate the effect of competing interactions on the structural heterogeneity. Our theory can be straightforwardly generalized to brushes with various chain architectures, copolymer compositions and charge patterns. It can also be easily implemented to study protein brushes with specific amino acid sequences. Similar multi-layer morphologies with separate chain populations in each layer have been found in our previous studies on intrinsically disordered protein brushes 69, 70.
Here, we focus on the multi-layer morphologies formed by microphase segregation of PE brushes in the normal direction. However, solving SCFT in high dimensions for brushes with varying grafting densities may lead to segregation in other directions with rich surface patterns 23. We leave the exploration of the full phase diagram for a future work. Furthermore, current applications of PE brushes largely rely on their steric and electrostatic repulsion to protect the surface underneath. As revealed in this work, designing structurally heterogeneous PE brushes by tuning their electrostatic and hydrophobic interactions could regulate surface properties at a higher level.
@@
Effect of bulk salt concentration on the brush morphology; electric double layer structure within multi-layer morphologies; effect of polymer dielectric constant on the stability of multilayer morphologies; effect of bulk salt concentration on the reflectivity spectra.
Supplemental Information
0.2 I. Effect of Bulk Salt Concentration on the Brush Morphology
We examine the dependence of the brush morphology on the bulk salt concentration. We plot the polymer density profile in Fig. S1 and brush height in Fig. S1. We find brushes shrink as salt concentration increases. The height almost follows the scaling of $H\sim(c_{\pm}^{b})^{-1/3}$, suggesting that the brushes studied here are in the salted regime. Therefore, adding salt indeed has an impact on the multi-layer morphology. Increasing the salt concentration enhances the screening effect and thus reduces the electrostatic repulsion within the brushes, sharing a similar effect as reducing the backbone charge density $\alpha$. We expect that as the salt concentration increases, the multi-layer structures become less accessible.
Figure S1: Effect of bulk salt concentration on (a) the density profile and (b) brush height $H$ of swollen PE brushes with $\alpha=0.4$ and $\chi=1.00$. The dotted line in (b) represents the scaling result for the salted regime, $H\sim(c_{\pm}^{b})^{-1/3}$.
[…]
0.3 II. Electric Double Layer Structure within Multi-layer Morphologies
Figure S2: Distributions of cations (dotted lines) and anions (dashed lines) in different multi-layer morphologies. The concentration of ions $c_{\pm}$ (normalized by the corresponding bulk concentration $c_{\pm}^{b}$) is plotted in the $z$-direction normal to the substrate. The parameters used in (a) to (f) correspond to those in Fig. 2(a) to (f) in the main text.
0.4 III. Effect of Polymer Dielectric Constant on the Stability of Multi-layer Morphologies
Our theory can also include the non-uniform dielectric permittivity. Here, we performed calculations using polymers with lower dielectric constant $\epsilon_{p}=10$. The local dielectric constant is taken to be the linear combination of that of the polymer and the solvent, $\epsilon_{r}(\mathbf{r})=\epsilon_{p}\phi_{p}(\mathbf{r})+\epsilon_{s}\phi_{s}(\mathbf{r})$. Fig. S3 shows the free energy as a function of $\chi$ for the multi-layer morphologies calculated at $\epsilon_{p}=10$. Compared to Fig. 4 in the main text, the transition points between consecutive layers shift to lower values of $\chi$. In general, lowering $\epsilon_{p}$ has the same effect as lowering the backbone charge density $\alpha$. The reason is that lowering $\epsilon_{p}$ increases the screening effect on the electrostatic repulsions within the layers, which makes the hydrophobic attraction more dominant in the competition. This effect suppresses multi-layer formation.
Figure S3: Free energy (in excess to the reference bulk salt solution) as a function of the hydrophobicity $\chi$ for $\epsilon_{p}=10$ and $\epsilon_{s}=80$ with fixed $\alpha=0.4$. The dotted lines indicate the metastable regimes. Vertical dashed lines indicate the locations of discontinuous transitions between morphologies. Inset shows excess free energies with respect to the morphology with two condensed layers to highlight the energy crossings. The local dielectric constant is calculated from the linear combination of that of the polymer and solvent, $\epsilon_{r}(\mathbf{r})=\epsilon_{p}\phi_{p}(\mathbf{r})+\epsilon_{s}\phi_{s}(\mathbf{r})$.
0.5 IV. Effect of Bulk Salt Concentration on the Reflectivity Spectra
We examine the salt effect on the spectra. Fig. S4 shows the effect of salt concentration on the density profile and reflectivity spectra of the two-layered morphology. Increasing the salt concentration increases the screening effect and thus makes the layers more compact. In the spectra, the peaks shift to higher $Q_{z}$.
Figure S4: Evolution of reflectivity spectra at three different salt concentrations for PE brushes with $\alpha=0.4$ and $\chi=1.45$. (a) Density distributions and (b) corresponding reflectivity spectra for PE brushes with two condensed layers. For clarity, the curves are offset vertically by 10 equiv as $c_{\pm}^{b}$ increases. The dotted guideline illustrates the shift of the spectra to higher $Q_{z}$.